OpenStudy (anonymous):

Find the local maximum and minimum values and saddle point(s) of the function f(x,y) = e^x cos(y)

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

I worked out fx and fy so\[f_x=e^x\cos(y)\text{ }f_y=-e^x\sin(y)\] So now?

OpenStudy (amistre64):

fxx fyy - (fxy)^2 rings a bell

OpenStudy (anonymous):

Yes, I have to determine that as well. But first I need to find the critical points, right? I'm a little lost... hah

OpenStudy (amistre64):

critical points are equal to 0 or undefined

OpenStudy (amistre64):

fx is just ignoring the y direction and focusing on how it curves along x

OpenStudy (anonymous):

Does it work like that every single time? I.e. when you look at fx=0, you ignore all the y's if there are any in the fx? Same for the fy?

OpenStudy (amistre64):

f = e^x cos(y) fx = e^x cos(y) fy = -e^x sin(y) fx = 0 when e^x = -cos(y) by ignoring the ys, i mean that y is simply some constant during a derivative

OpenStudy (amistre64):

fxx tells us curvative

OpenStudy (amistre64):

fxx = e^x cos(y) fyy = -e^x cos(y) fxy = -e^x sin(y)

OpenStudy (anonymous):

Oh haha! so how do I get the values when I say fx and fy = 0? Im struggling with them in problems like this one

OpenStudy (amistre64):

i think we should get it to this point, then compare the rules: D = fxx fyy - (fxy)^2 if D>0 and fxx > 0, then min at a,b if D>0 and fxx < 0, then max at a,b if D<0, then saddle point if D=0, then its still undetermined

OpenStudy (anonymous):

Okay

OpenStudy (amistre64):

also, there is something about a system of equations for fx fy fx = 0 fy = 0 solve the system of equations

OpenStudy (amistre64):

e^x cos(y) = -e^x sin(y) cos(y) = -sin(y) ; this is a 45 degree angle in Q2 and Q4 if my minds eye is not blind

OpenStudy (anonymous):

Okay so e^x cos(y) = 0 -e^(x) sin(y) = 0 How to I get the values of x and y? In fx we know that either/both of e^x or cos y makes the answer be 0. It cannot be e^x, right? So we are left with cos y = 0, right?

OpenStudy (anonymous):

Oh I see! I can say 0 = -sin(y)/cos(y) = -tan(y) and solve for y plug that in to any eq to get x?

OpenStudy (amistre64):

sounds like a winner to me

OpenStudy (anonymous):

Was my first attempt wrong?

OpenStudy (amistre64):

i would say -1 = tan(y) is a solution, not 0=tan(y)

OpenStudy (amistre64):

y = 3pi/4 or -pi/4

OpenStudy (anonymous):

I didn't get that with -tan(y) = 0. I got y = 0

OpenStudy (anonymous):

Ah I see, I made a stupid mistake haha

OpenStudy (amistre64):

thats because its not 0 :) e^x cos(y) = -e^x sin(y) -cos(y) = sin(y) -1 = tan(y) sin(-pi/4) = -sqrt(2)/2 cos(-pi/4) = sqrt(2)/2 sin(3pi/4) = sqrt(2)/2 cos(3pi/4) = -sqrt(2)/2

OpenStudy (anonymous):

Thanks @amistre64 ! I got the correct answer now!

OpenStudy (amistre64):

yay!! :)